# Avoiding Catastrophe in Online Learning by Asking for Help

### v4

@@ Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
This section presents another algorithm with subconstant regret and sublinear queries, but under different assumptions. The primary takeaway here is that our algorithm can handle the class of thresholds on $[0,1]$, which is known to have infinite Littlestone dimension and thus be hard in the standard online learning model. (Example 21.4 in Shalev-Shwartz & Ben-David (2014)).
@@
Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actionx) on $\boldsymbol{x}$. For example, if $\Pi$ is the class of thresholds on $[0,1]$, every $\pi\in\Pi$ has at most two segments, and thus the positive result in this section will apply. This demonstrates the existence of policy classes which are learnable in our setting but not learnable in the standard online learning model, meaning that the two settings do not exactly coincide.
We prove the following regret bound. Like our other positive results, this bound applies to both multiplicative and additive regret.
@@
For any $\boldsymbol{x}\in\mathcal{X}^{T}$, any $\pi^{m}$ with $S(\pi^{m})\leq K$, and any function $g:\mathbb{N}\to\mathbb{N}$, Algorithm 4 makes at most $(\operatorname*{diam}(\boldsymbol{x})+4)g(T)$ queries and satisfies
$R_{T}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ \frac{2LKT}{g(T)^{2}}$
@@
For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ makes $O(T^{c}(\operatorname*{diam}(\boldsymbol{x})+1))$ queries and satisfies
$R_{T}\leq$ $\ 2LKT^{1-2c}$
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ 2LKT^{1-2c}$
Our algorithm does not need to know $L$ or the number of segments; it only needs to know $T$.
@@ D.1 Intuition behind the algorithm
The algorithm maintains a set of buckets which partition the observed portion of the input space. Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the input space. As long as the bucket contains a query from a prior time step, local generalization allows us to bound $\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})$ based on the length of the bucket containing $x_{t}$. We always query if the bucket does not contain a prior query
@@
1:function AvoidCatastrophe($T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N}$)
2:     $X_{Q}\leftarrow\emptyset$triangleright Previously queried inputs
3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{m}(x)$ for each $x\in X_{Q}$
4:     $\mathcal{B}\leftarrow\emptyset$triangleright The set of active buckets
5:     for $t$ from $1$ to $T$ do
6:         EvaluateInput$(x_{t})$
7:function EvaluateInput($x\in\mathcal{X}$)
8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $x$: create a new bucket and try again
9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $x\in B$
10:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
11:         $n_{B}\leftarrow 0$triangleright Number of time steps that have used $B$
12:         EvaluateInput$(x)$
13:     else
14:         $B\leftarrow$ any bucket containing $x$
15:         if $X_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
16:              Query mentor and observe $\pi^{m}(x)$
17:              $\pi(x)\leftarrow\pi^{m}(x)$
18:              $X_{Q}\leftarrow X_{Q}\cup\{x\}$
19:              $n_{B}\leftarrow n_{B}+1$
20:         else if $n_{B}<T/g(T)$ then $\triangleright$ Bucket has a query and isn’t full: take that action
21:              Let $x^{\prime}\in X_{Q}\cap B$
22:              Take action $\pi(x^{\prime})$
23:              $n_{B}\leftarrow n_{B}+1$
24:         $\triangleright$ Bucket is full: split bucket and try again
25:              $B=[a,b]$
26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[a,\frac{a+b}{2}\right],\left[\frac{a+b}{2},b\right]\Big{)}$
27:              $(x_{B_{1}},x_{B_{2}})\leftarrow(0,0)$
28:              $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
29:              EvaluateInput$(x)$
D.2 Notation for the proof
@@
    Let $B_{t}$ be the bucket that is used on time step $t$ (as defined on line 14 of Algorithm 4).
@@
      Buckets created on line 9 are depth 0.
@@
      We refer to $B_{1},B_{2}$ created on line 26 as the children of the bucket $B$ defined on line 14.
@@
    Viewing the set of buckets are a binary tree defined by the “child” relation, we use the terms “ancestor” and ”descendant” in accordance with their standard tree definitions.
@@ D.3 Proof roadmap
    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma D.4), with Lemma D.3 as an intermediate step. This captures the total amount of suboptimality.
@@
    As in the proof of Theorem 5.2, Lemma B.3 transforms the multiplicative objective into an additive form. Lemma D.5 bounds the additive objective using Lemmas D.2 and D.4.
@@
    Combining Lemmas D.5 and B.3 bounds the regret (Lemma D.6).
  6. 6.
    Theorem D.2 directly follows from Lemmas D.1 and D.6.
@@ D.4 Proof
Algorithm 4 performs at most one query per bucket, so the total number of queries is bounded by the total number of buckets. There are two ways to create a bucket: from scratch (line 9), or by splitting an existing bucket (line 26).
@@
If we query the mentor at time $t$, $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$. Thus assume we do not query the mentor at time $t$: then there exists $x^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$. Since $x_{t}$ and $x^{\prime}$ are both in $B_{t}$, $|x_{t}-x^{\prime}|\leq\operatorname*{len}(B_{t})$. Then by the local generalization assumption, $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$. ∎
@@
Next, observe that every $B\in\mathcal{B}_{V}\setminus\mathcal{B}_{V}^{\prime}$ must have a descendent in $\mathcal{B}_{V}^{\prime}$: otherwise we would have $B\in\mathcal{B}_{V}^{\prime}$. Let $\mathcal{A}(B)$ denote the set of ancestors of $B$, plus $B$ itself. Then we can write
@@
$\sum_{B\in\mathcal{A}(B^{\prime})}$ |$\ \frac{|\{t\in M_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq\frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
$=$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{d(B^{\prime})}\frac{1}{2^{d}}\leq\frac{T}{g(T)^{2}}\sum_{d=0}^{\infty}\frac{1}{2^{d}}=\frac{2T}{g(T)^{2}}$
@@
For $t\not\in M_{T}$ we have $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=\mu_{t}^{m}(x_{t})$ by definition, and Lemma D.2 implies that $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\geq L\operatorname*{len}(B_{t})$ for all $t\in[T]$. Thus
$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq$ $\ \sum_{t\in M_{T}}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
$\leq$ $\ L\sum_{t\in M_{T}}\operatorname*{len}(B_{t})$
Then by Lemma D.4,
$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
as required. ∎
Lemma D.6.
Under the conditions of Theorem D.2, Algorithm 4 satisfies
$R_{T}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ \frac{2LKT}{g(T)^{2}}$
Proof.
Let $a_{t}=\mu_{t}^{m}(x_{t})$ and $b_{t}=\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
$\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
Since $\mu_{t}(x_{t},y_{t})\geq\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$, we have
$R_{T}=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
Similarly,
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$
Applying Lemma D.5 to the above two inequalities completes the proof. ∎
@@
Appendix E Other proofs
@@
Thus $-2L||x-x^{\prime}||\leq\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)\leq 2L||x-x^{\prime}||$. This is equivalent to $|\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)|\leq 2L||x-x^{\prime}||$, completing the proof.
Proposition E.2 states that whether subconstant regret is achievable does not depend on whether we require expected sublinear queries or worst-case sublinear queries.
Proposition E.2.
Suppose an algorithm satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\in o(T)$. Then there exists $h:\mathbb{N}\to\mathbb{N}$ such that (1) $h(T)\in o(T)$ and (2) if the algorithm is modified to simply stop querying if the number of queries reaches $h(T)$, the algorithm still satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
@@
We use $Q_{T},R_{T}$ to refer to the queries and regret of the original algorithm, and $Q_{T}^{\prime},R_{T}^{\prime}$ to refer to the queries and regret of the modified algorithm.
Since $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\in o(T)$, there exists $g:\mathbb{N}\to\mathbb{N}$ such that $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$ and $g(T)\in o(T)$. Let $h(T)=\sqrt{g(T)T}$; then $h(T)\in o(T)$ by Lemma A.1. Markov’s inequality implies that
$\Pr\big{[}|Q_{T}|>h(T)\big{]}\leq\frac{\operatorname*{\mathbb{E}}[|Q_{T}|]}{h(T)}\leq\frac{g(T)}{\sqrt{g(T)T}}=\sqrt{\frac{g(T)}{T}}$
Let $\xi$ denote the event that at some point, the original algorithm would query, but the modified algorithm cannot because $|Q_{T}^{\prime}|=h(T)$. Then $\Pr[\xi]\leq\Pr[|Q_{T}|>h(T)]$ (the inequality is because the modified algorithm might not want to query more anyway). Also note that the algorithms are equivalent if $\xi$ does not occur, so $\operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\neg\xi]=\operatorname*{\mathbb{E}}[R_{T}]$. Hence
$\operatorname*{\mathbb{E}}[R_{T}^{\prime}]=$ $\ \operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\neg\xi]\Pr[\neg\xi]+\operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\xi]\Pr[\xi]$
$\leq$ $\operatorname*{\mathbb{E}}[R_{T}]\cdot 1+1\cdot\Pr[\xi]$
$\leq$ $\ \operatorname*{\mathbb{E}}[R_{T}]+\sqrt{\frac{g(T)}{T}}$
Since $g(T)\in o(T)$, we get
$\lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}[R_{T}^{\prime}]\leq$ $\ \lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \left(\operatorname*{\mathbb{E}}[R_{T}]+\sqrt{\frac{g(T)}{T}}\right)$
$=$ $\ \lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}[R_{T}]+\lim_{T\to\infty}\ \sqrt{\frac{g(T)}{T}}$
[…]
as required. ∎
Like the proof of Theorem 4.1, the proof of Theorem E.3 assumes worst-case sublinear queries. However, recall from Proposition E.2 that the distinction between worst-case sublinear queries and expected sublinear queries is not significant.
Theorem E.3.
Let $\mathcal{X}=[0,1]$ and $\mathcal{Y}=\{0,1\}$. Assume each input is sampled i.i.d. from the uniform distribution on $\mathcal{X}$ and define the mentor policy class by the set of intervals within $\mathcal{X}$, i.e., $\Pi=\{\pi:\exists a,b\in[0,1]\text{ s.t }\pi(x)=\mathbf{1}(x\in[a,b])\ \forall x\in\mathcal{X}\}$. Then without the local generalization assumption, any algorithm with sublinear queries satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=1$.
Proof.
Part 1: Setup. Consider any algorithm which makes sublinear worst-case queries: then there exists $g:\mathbb{N}\to\mathbb{N}$ where $\sup_{\boldsymbol{\mu},\pi^{m}}|Q_{T}|\leq g(T)$ and $g(T)\in o(T)$. Define $f(T):=\sqrt{(g(T)+1)T}$; by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Divide $\mathcal{X}$ into $f(T)$ equally sized sections $X_{1},\dots,X_{f(T)}$ in the exactly the same way as in Section 4.2; see also Figure 2. Assume that each $x_{t}$ is in exactly one section: this assumption holds with probability 1, so it does not affect the regret.
We use the probabilistic method: sample a segment $j^{m}\in[f(T)]$ uniformly at random, define $\pi^{m}$ by $\pi^{m}(x)=\mathbf{1}(x\in X_{j^{m}})$, and define $\mu$ by $\mu(x,y)=\mathbf{1}(y=\pi^{m}(x))$. In words, the mentor takes action 1 iff the input is in section $j^{m}$, and the agent receives payoff 1 if its action matches the mentor’s and zero otherwise. Since any choice of $j^{m}$ defines a valid $\mu$ and $\pi^{m}$, we have
$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections which are never queried. Let $j_{1},\dots,j_{k}$ be the sequence of sections queried by the agent: then $k\leq|Q_{T}|\leq g(T)$.
@@
Now fix $i$ and assume $j_{r}\neq j^{m}\ \forall r<i$. Queries in sections other than $j^{m}$ provide no information about the value of $j^{m}$, so $j^{m}$ is uniformly distributed across the set of sections not yet queried, i.e., $\{j\in[f(T)]:j_{r}\neq j\ \forall r<i\}$. There are at least $f(T)-i+1$ such sections, since there are $i-1$ prior queries at this point. Thus $\Pr[j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i]\leq\frac{f(T)-i}{f(T)-i+1}$ (the inequality is because it could also be 0 if $j_{i}=j_{r}$ for some $i<r$). Therefore
$\Pr\big{[}j^{m}\in J_{\neg Q}\big{]}\leq$ $\ \prod_{i=1}^{k}\frac{f(T)-i}{f(T)-i+1}$

### v5

@@ Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
This section presents another algorithm with subconstant regret and sublinear queries, but under different assumptions. The primary takeaway here is that our algorithm can handle the class of thresholds on $[0,1]$, which is known to have infinite Littlestone dimension and thus be hard in the standard online learning model. (Example 21.4 in Shalev-Shwartz & Ben-David, 2014).
@@
Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actions) on $\boldsymbol{x}$. For example, if $\Pi$ is the class of thresholds on $[0,1]$, every $\pi\in\Pi$ has at most two segments, and thus the positive result in this section will apply. This demonstrates the existence of policy classes which are learnable in our setting but not learnable in the standard online learning model, meaning that the two settings do not exactly coincide.
Unlike our primary algorithm (Algorithm 1), this algorithm does require direct access to the input encoding. However, the point of this section is not to present a practical algorithm: it is simply to demonstrate that our setting and the standard online setting do not exactly coincide.
We prove the following regret bound. Like our previous results, this bound applies to both multiplicative and additive regret.
@@
For any $\boldsymbol{x}\in\mathcal{X}^{T}$, any $\pi^{m}$ with $S(\pi^{m})\leq K$, and any function $g:\mathbb{N}\to\mathbb{N}$ satisfying $g(T)\geq 2L/\mu_{0}^{m}$, Algorithm 4 satisfies
$R_{T}^{\times}\leq$ $\ \frac{4LKT}{g(T)^{2}\mu_{0}^{m}}$
$R_{T}^{+}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
$|Q_{T}|\leq$ $\ (\operatorname*{diam}(\boldsymbol{x})+4)g(T)$
@@
For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ satisfies
$R_{T}^{\times}\in$ $\ O\left(\frac{LKT^{1-2c}}{\mu_{0}^{m}}\right)$
$R_{T}^{+}\in$ $\ O\left(LKT^{1-2c}\right)$
$|Q_{T}|\in$ $\ O(T^{c}(\operatorname*{diam}(\boldsymbol{x})+1))$
1:  function DBWRQ$(T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N})$
2:     $X_{Q}\leftarrow\emptyset$  (previously queried inputs)
3:     $\pi\leftarrow\emptyset$  (records $\pi^{m}(x)$ for each $x\in X_{Q}$)
4:     $\mathcal{B}\leftarrow\emptyset$  (The set of active buckets)
5:     for $t$ from $1$ to $T$ do
6:        EvaluateInput$(x_{t})$
7:     end for
8:  end function
9:  function EvaluateInput$(x\in\mathcal{X})$
10:     if $x\not\in B$ for all $B\in\mathcal{B}$ then
11:        $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $x\in B$
12:        $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
13:        $n_{B}\leftarrow 0$
14:        EvaluateInput$(x)$
15:     else
16:        $B\leftarrow$ any bucket containing $x$
17:        if $X_{Q}\cap B=\emptyset$ then
18:           Query mentor and observe $\pi^{m}(x)$
19:           $\pi(x)\leftarrow\pi^{m}(x)$
20:           $X_{Q}\leftarrow X_{Q}\cup\{x\}$
21:           $n_{B}\leftarrow n_{B}+1$
22:        else if $n_{B}<T/g(T)$ then
23:           Let $x^{\prime}\in X_{Q}\cap B$
24:           Take action $\pi(x^{\prime})$
25:           $n_{B}\leftarrow n_{B}+1$
26:        else
27:           $B=[a,b]$
28:           $(B_{1},B_{2})\leftarrow\Big{(}\left[a,\frac{a+b}{2}\right],\left[\frac{a+b}{2},b\right]\Big{)}$
29:           $(x_{B_{1}},x_{B_{2}})\leftarrow(0,0)$
30:           $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
31:           EvaluateInput$(x)$
32:        end if
33:     end if
34:  end function
@@ D.1 Intuition behind the algorithm
We call our algorithm “Dynamic Bucketing With Routine Querying”, or DBWRQ (pronounced “DBWRQ”). The algorithm maintains a set of buckets which partition the observed portion of the input space. Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the input space. As long as the bucket contains a query from a prior time step, local generalization allows us to bound $\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})$ based on the length of the bucket containing $x_{t}$. We always query if the bucket does not contain a prior query; in this sense the querying is “routine”.
@@
D.2 Proof notation
@@
    Let $B_{t}$ be the bucket that is used on time step $t$ (as defined on line 16 of Algorithm 4).
@@
      Buckets created on line 11 are depth 0.
@@
      We refer to $B_{1},B_{2}$ created on line 28 as the children of the bucket $B$ defined on line 16.
@@
    Viewing the set of buckets as a binary tree defined by the “child” relation, we use the terms “ancestor” and ”descendant” in accordance with their standard tree definitions.
@@ D.3 Proof roadmap
    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma D.4), with Lemma D.3 as an intermediate step. This captures the “total amount of suboptimality”.
@@
    Lemma D.5 uses Lemma D.2 and Lemma D.4 to bound the regret.
@@
    Theorem D.2 directly follows from Lemmas D.1 and D.5.
@@ D.4 Proof
Algorithm 4 performs at most one query per bucket, so the total number of queries is bounded by the total number of buckets. There are two ways to create a bucket: from scratch (line 11), or by splitting an existing bucket (line 28).
@@
If we query at time $t$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$. Thus assume we do not query at time $t$: then there exists $x^{\prime}\in B_{t}$ (as defined on line 23 of Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$. Since $x_{t}$ and $x^{\prime}$ are both in $B_{t}$, $|x_{t}-x^{\prime}|\leq\operatorname*{len}(B_{t})$. Then by local generalization, $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$. ∎
@@
Next, observe that every $B\in\mathcal{B}_{V}\setminus\mathcal{B}_{V}^{\prime}$ must have a descendant in $\mathcal{B}_{V}^{\prime}$: otherwise we would have $B\in\mathcal{B}_{V}^{\prime}$. Let $\mathcal{A}(B)$ denote the set of ancestors of $B$, plus $B$ itself. Then we can write
@@
$\sum_{B\in\mathcal{A}(B^{\prime})}\frac{|\{t\in M_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq$ $\ \frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
$=$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{d(B^{\prime})}\frac{1}{2^{d}}$
$\leq$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{\infty}\frac{1}{2^{d}}$
$=$ $\ \frac{2T}{g(T)^{2}}$
@@
We have
[…]
Since $g(T)\geq 2L/\mu_{0}^{m}$ and every bucket length is at most $\frac{1}{g(T)}$,
$\mu_{t}(x_{t},y_{t})\geq$ $\ \mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$
$\geq$ $\ \mu_{0}^{m}-\frac{L}{g(T)}$
$\geq$ $\ \mu_{0}^{m}-\mu_{0}^{m}/2$
$\geq$ $\ \mu_{0}^{m}/2$
[…]
Invoking Lemma A.1 completes the proof:
$R_{T}^{\times}\leq\frac{2R_{T}^{+}}{\mu_{0}^{m}}\leq\frac{4LKT}{g(T)^{2}\mu_{0}^{m}}$
@@
Appendix E Properties of local generalization
@@
So $-2L||x-x^{\prime}||\leq\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)\leq 2L||x-x^{\prime}||$ and therefore $|\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)|\leq 2L||x-x^{\prime}||$.∎
Note that Theorem E.2 as stated only provides a bound on $R_{T}^{\times}$. A similar bound can be obtained for $R_{T}^{+}$, but it is more tedious and we do not believe it would add much to the paper.
Theorem E.2.
Let $\mathcal{X}=[0,1]$ and $\mathcal{Y}=\{0,1\}$. Let each input be sampled i.i.d. from the uniform distribution on $\mathcal{X}$ and define the mentor policy class as the set of intervals within $\mathcal{X}$, i.e., $\Pi=\{\pi:\exists a,b\in[0,1]\text{ s.t }\pi(x)=\mathbf{1}(x\in[a,b])\ \forall x\in\mathcal{X}\}$. Then without the local generalization assumption, any algorithm with sublinear queries satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{\times}]=\infty$.
@@
Part 1: Setup. Consider any algorithm which makes sublinear worst-case queries: then there exists $g:\mathbb{N}\to\mathbb{N}$ where $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$ and $g(T)\in o(T)$. WLOG assume $g(T)\geq 0$ for all $T$; if not, redefine $g(T)$ to be $\max(g(T),1)$.
Define $f(T):=\lceil\sqrt{g(T)T}\rceil$; by Lemma A.2, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Divide $\mathcal{X}$ into $f(T)$ equally sized sections $X_{1},\dots,X_{f(T)}$ in the exactly the same way as in Section 4.2; see also Figure 1. Assume that each $x_{t}$ is in exactly one section: this assumption holds with probability 1, so it does not affect the regret.
We use the probabilistic method: sample a segment $j^{m}\in[f(T)]$ uniformly at random, define $\pi^{m}$ by $\pi^{m}(x)=\mathbf{1}(x\in X_{j^{m}})$, and define $\mu$ by $\mu(x,y)=\mathbf{1}(y=\pi^{m}(x))$. In words, the mentor takes action 1 iff the input is in section $j^{m}$, and the agent receives payoff 1 if its action matches the mentor’s and zero otherwise. Since any choice of $j^{m}$ defines a valid $\mu$ and $\pi^{m}$,
$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections which are never queried. Let $j_{1},\dots,j_{k}$ be the sequence of sections queried by the agent: then $k=|Q_{T}|\leq g(T)$.
@@
Now fix $i$ and assume $j_{r}\neq j^{m}\ \forall r<i$. Queries in sections other than $j^{m}$ provide no information about the value of $j^{m}$, so $j^{m}$ is uniformly distributed across the set of sections not yet queried, i.e., $\{j\in[f(T)]:j_{r}\neq j\ \forall r<i\}$. There are at least $f(T)-i+1$ such sections, since there are $i-1$ prior queries at this point. Thus $\Pr[j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i]\geq\frac{f(T)-i}{f(T)-i+1}$ (the inequality is because this probability could also be 1 if $j_{i}=j_{r}$ for some $i<r$). Therefore
$\Pr\big{[}j^{m}\in J_{\neg Q}\big{]}\geq$ $\ \prod_{i=1}^{k}\frac{f(T)-i}{f(T)-i+1}$